OpenStudy (anonymous):

Find the area of the triangle with a = 11.8, b = 12.6, c = 14.8. Round to the nearest tenth.

OpenStudy (amistre64):

sounds like a herons rule ....

OpenStudy (austinl):

I was thinking about that myself.

OpenStudy (amistre64):

i hate herons rule :)

OpenStudy (austinl):

Nah, it is simple enough.

OpenStudy (austinl):

It is a two step process, 1) We need to calculate "s", which is half of the triangle perimeter. \(\displaystyle s=\frac{a+b+c}{2}\) 2) We can then calculate the area using a neat little formula! \(A=\sqrt{s(s-a)(s-b)(s-c)}\)

OpenStudy (amistre64):



OpenStudy (amistre64):

lol, i swear i can find the hardest way to do the simplest tasks :)

OpenStudy (austinl):

It is a gift and a curse :P

OpenStudy (anonymous):

okay, I'm working it out now. Thanks guys

OpenStudy (amistre64):

i developed a right triangle with the given ab, and scaled it to the given c is all

OpenStudy (amistre64):

n = 14.8/sqrt(298) area = 12.6n * 11.8/2

OpenStudy (anonymous):

I got 71.7